The square of a number decreased by 2 is equal to the negative of the number. Find the number.
Answer by ankor@dixie-net.com(12701)

The square of a number decreased by 2 is equal to the negative of the number. Find the number.
:
Let "a number" = x
Squared: x^2
decreased by 2: (x^2 - 2)
negative of the number: -x
:
x^2 - 2 = -x
:
add x to both sides, you then have a quadratic equation:
x^2 + x - 2 = 0
:
Factors to:
(x + 2)(x - 1) = 0
:
x = -2;
and
x = +1
:
Check using x = -2 in the original equation:
(-2)^2 + (-2) = -(-2)
+4 - 2 = +2
:
Check using x = +1
(1^2) - 2 = -1
1 - 2 = -1